OpenStudy (anonymous):

Michiko works in a shop that rents cross-country skis. Two clients come in to rent skis. Michiko knows that a person's height is proportional to his or her correct ski length. She figures out that the first client, who is 70 inches tall, needs 210-centimeter skis. The second client is 63 inches tall. What length of skis does the second client need?

OpenStudy (anonymous):

Solve for x.\[\frac{70}{210}=\frac{63}{x} \]

OpenStudy (anonymous):

70-210 63-X 210*63/70=189 X=189 cemtimeters